OpenStudy (anonymous):

A travel mug of 90∘C coffee is left on the roof of a parked car on a cold winter day. The temperature of the coffee t minutes after it was left on the roof is given by H(t)=90(2)−t/13. After how many minutes will the coffee be only lukewarm (30∘C)?

OpenStudy (anonymous):

\[-t/13=\log_{2} (30/90)\]

OpenStudy (anonymous):

This is as far as I got ^^

OpenStudy (campbell_st):

well you know h(t) = 30 so make the substitution so \[30 = 90 \times (2)^{-\frac{t}{13}}\] divide both sides by 90 \[\frac{1}{3} = (2)^{-\frac{t}{13}}\] now take the log of both sides of the equation, using log laws for the power \[\ln(\frac{1}{3}) = -\frac{t}{13} \times \ln(2)\] hopefully you can now solve for t

OpenStudy (campbell_st):

as a tip, don't use base 2 logs... use a log that is easy to calculate with, natural logs or base 10 logs... both will give the same answer.

OpenStudy (empty):

Yeah if you're good you can leave out the base of your logarithm until you're ready to evaluate it.